OpenStudy (anonymous):

greatest common factor of 270 and 360

OpenStudy (anonymous):

270 = 27*10 = 2*3^3*5 360 = 10*36 = 2*5*(2*3)^2 = 2^3*3^2*5

OpenStudy (anonymous):

now just pick the lowest exponent in each common factor: 2*3^2*5